OpenStudy (bimbim2697):

SOLVE EACH SYSTEM OF EQUATIONS BY ELIMINATION.TELL WHETER THE SYSTEM HAS ONE SOLUTION INFINITE SOLUTION MANY SOLUTIONS OR NO SOLUTION.CHECK YOUR SOLUTION 1)4X+Y=8 X+Y=5 2)3X-2Y=10 X-2Y=6

OpenStudy (radar):

1. has one solution, x=1, y=4 2. has one solution, x=2, y=-2

OpenStudy (bimbim2697):

@radar can u help me with another question

OpenStudy (bimbim2697):

A farm raises a total of 220 chickens and pigs.The number of legs of the stock in the farm totals 520 .How many chickens and pigs are at the farm?

OpenStudy (radar):

Number of animals is 220 given. Let p = pigs Let c = chickens. Then c + p =220 Now we do the legs. Here is a fundamental fact, chickens have two legs and pigs have 4 legs. We now can write: 2c + 4p = 520 Substitute for c, in that c=220-p 2(220-p) + 4p = 520 440 -2p +4p = 520 2p =520 -440 2p = 80 p = 40 forty pigs c = 180 chickens